OpenStudy (anonymous):

Determine S200 for each of the following arithmetic series 3+6+9+12+...

OpenStudy (nenadmatematika):

the first element is 3 and an=a1+(n-1)*d so an=3+199*3=600 so the sum is (200/2)*(3+600)=100*603=60300

OpenStudy (anonymous):

a (200th) = 3 + 199 * 3 = 600 S ( 200) = 200 ( 3 + 600)/2 = 60300